OpenStudy (anonymous):

(1/4)X + 24 = 36 12 4 48 24

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }x+24=36\] \[\frac{ 1 }{ 4 }x=36-24\] \[\frac{ 1 }{ 4 }x=12\] then, times both sides with 4

OpenStudy (anonymous):

x=48 :)